you invested $25,000 in two accounts paying 6% and 8% interest, respectively. if the total interest earned for the year was $1780, how much was invested at each rate?
Answer by mananth(16946)

Part I 6.00% per annum ------------- Amount invested =x
Part II 8.00% per annum ------------ Amount invested = y
25000
Interest----- 1780.00

Part I 6.00% per annum ---x
Part II 8.00% per annum ---y
Total investment
x + 1 y= 25000 -------------1
Interest on both investments
6.00% x + 8.00% y= 1780
Multiply by 100
6 x + 8 y= 178000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -150000.00
Add this to (2)
0 x 2 y= 28000
divide by 2
y = 14000 - x)
Part I 6.00% $ 11000
Part II 8.00% $ 14000

CHECK
11000 --------- 6.00% ------- 660.00
14000 ------------- 8.00% ------- 1120.00
Total -------------------- 1780.00

m.ananth@hotmail.ca